. Select the most polar bond amongst the following. A) C-O B) Si-F C) CI-F D) C-F E) C-I 2. Hydrazine, N2H, is a good reducing agent that has been used as a component in rocket fuels. Select its Lewis structure A) B) C) D H-N-N-H E) None of the above is correct. 3. Select the best Lewis structure for CICN D) CI-C-N E) :Cic-N
Answer #1

1. The electronegativity difference between the atoms of the most polar bond is the highest.

A. C-O:

The electronegativity value of C is 2.5.

The electronegativity value of O is 3.5.

Difference = 3.5-2.5 = 1

B. Si-F:

The electronegative value of Si is 1.8.

The electronegativity value of F is 4

Difference = 4-1.8 = 2.2

C. Cl-F:

The electronegativity value of Cl is 3.

The electronegativity value of F is 4.

Difference = 1.

D. C-F:

Differene = 4-2.5 = 1.5

E. C-I

Difference = 2.5-2.5 = 0

Therefore, option B. Si-F is most polar bond.